OpenStudy (anonymous):

What is the minimum 3x^2+2^6 x^-3

OpenStudy (anonymous):

I believe it is 20 when x=2. (2, 20). Try graphing it if you are allowed to.

OpenStudy (anonymous):

How did you do that?

OpenStudy (raden):

fisrt, can you find the derivative of 3x^2+2^6 x^-3 ?

OpenStudy (anonymous):

6x+12^5-3x^-4

OpenStudy (anonymous):

I found the derivative and checked it at critical points. When the derivative=0 or when it doesn't exist we have a critical point.

OpenStudy (raden):

12^5 (part) ? how you got that, @psam24 ?

OpenStudy (anonymous):

That critical point can then be used to determine if the point is a maximum or minimum. What we can do is set up a sort of number line known as a 'sign chart' in which we put critical value, and figure out if the points to the left and right are positive and negative.

OpenStudy (anonymous):

@RadEn thats the part im confused about what it the derivative of that, just 0?

OpenStudy (anonymous):

To find the derivative of 3x^2+2^6x^-3. First use the power rule on 3x^2. Then the quotient rule on 2^6x^-3.

OpenStudy (raden):

see the original function, let it is f(x) = 3x^2+2^6 x^-3 notice that 2^6 = 64, so f(x) = 3x^2+2^6 x^-3 = 3x^2+64 x^-3 therefore, f '(x)= 6x + 6(-3)x^-4 = 6x - 192x^-4

OpenStudy (anonymous):

Now that you have that derivative from Rad, set f'(x)=0 and see what you get.

OpenStudy (raden):

opss, i mean f '(x)= 6x + 64(-3)x^-4 = 6x - 192x^-4

OpenStudy (anonymous):

@RadEn ok lol i knew something was wrong then what?

OpenStudy (anonymous):

Please look at my above statement.

OpenStudy (anonymous):

what do you mean from your above statement, can i see it in action?

OpenStudy (anonymous):

Here...RadEn helped you get the derivative. f '(x)= 6x + 64(-3)x^-4 = 6x - 192x^-4 When f'(x)=0 or it doesn't exist, we can get something called a critical point. Critical points can help us figure out maximum and minimums. Set the derivative=0 and solve for x.

OpenStudy (anonymous):

wow ok that looks a bit tough

OpenStudy (anonymous):

Not particularly. 6x-192/x^4=0, start by telling me where this derivative cannot exist. (Hint, can you ever divide by 0?)

OpenStudy (anonymous):

how did you get to that next step of 6x-192/x^4=0

OpenStudy (anonymous):

Are you in calculus?

OpenStudy (anonymous):

We have an equation you gave us. f(x)=3x^2+2^6*x^-3 We have taken the derivative. f'(x)= 6x-192*x^-4 Now we set the derivative equal to 0. f'(x)=0, thus 6x-192*x^-4=0 x^-4 is the same as 1/x^4 so I can also say 6x-192/x^4=0

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

then from there?

OpenStudy (anonymous):

First tell me what you get for x so that the derivative is nonexistant. Then tell me what you get for x when 6x-192/x^4=0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Good! Now what do you get for x when 6x-192/x^4=0? Try factoring out an x.

OpenStudy (anonymous):

..

OpenStudy (raden):

f ' (x) = 0 6x - 192/x^4 = 0 6x = 192x^4 times x^4 on both sides, get 6x^5 = 192 divided by 6 both sides get x^5 = 192/6 x^5 = 32 setting 32 = 2^5, so x^5 = 2^5 the power both sides is same, therefore the bases must be same too, get x = 2 subtitute x = 2 into f(x) = 3x^2+64 x^-3 = 3x^2 + 64/ x^3 so, the value of f(2) = 3(2)^2 + 64/(2)^3 = 3(4) + 64/8 =12 + 8 = 20. that's the minimum value of f(x)